# REASSEMBLE: A Multimodal Dataset for Contact-rich Robotic Assembly and Disassembly

### v1

@@ III-D Data Collection
During teleoperation, all sensory information is recorded using the rosbag tool. Subsequently, the raw recordings are post-processed into HDF5 files, with the structure shown in Figure 2. We do not downsample or synchronize any sensory data during postprocessing; instead, we save the raw data along with timestamps. This approach allows for later adjustments in synchronization or data downsampling for specific downstream tasks. To optimize memory usage, all video and audio recordings are stored as encoded MP4 and MP3 files, respectively, which reduces the dataset size significantly, by a factor of six compared to the unprocessed rosbag recordings.
The accuracy of the predicted text for action narration depends on how clearly the operator spoke during data collection. We observed that some errors occur in the transcriptions; for example, ”D-SUB” is often recognized as ”the sub.” Such errors can be easily corrected automatically by substituting incorrect phrases with the correct ones. Therefore, we manually validate the data stored in the recordings and the action segment and success annotations. For this, we leverage the data visualization tool described in Section III-E. We verify that all data sources cover the entire recording, are synchronized, and that the action and success labels accurately reflect the events in the videos. If any labeling errors are identified, we correct them. Additionally, if any data is missing, we record the corresponding file name. Finally, based on the known sets of objects and actions, we can compute all valid annotations. We leverage this information to validate whether all action annotations are correct and ensure that no transcription mistakes remain in the dataset.
III-E Data Visualization
An overview of the data visualization tool is shown in the Appendix A.
@@ IV-A Action distribution
The number of demonstrations for each action-object pair and their distribution provide insights into the diversity and balance of the REASSEMBLE dataset. To achieve the best performance across all actions, it is crucial that the dataset contains a large and relatively equal number of demonstrations for each action. This ensures that downstream models observe each unique action an equal number of times, reducing performance bias toward more frequently occurring actions. The REASSEMBLE dataset includes four actions and 17 different objects, resulting in a total of 68 unique action-object pairs, or 69 when including the ”Idle” action. Figure 3 shows the number of occurrences for each unique action. For clarity, we present one plot per verb in the REASSEMBLE dataset.
The REASSEMBLE dataset contains a minimum of 55 demonstrations for the ”Remove square peg 2” action and a maximum of 86 demonstrations for the ”Pick USB” and ”Insert BNC” actions. This demonstrates that the dataset is balanced.
@@ IV-B Action difficulty and failure modes
The number of failed demonstrations per action can serve as a metric for task difficulty, as operators are more likely to fail when the motion is complex. Figure 4 shows the relationship between successful and unsuccessful demonstrations for each unique action. From the figure, we observe that the most difficult action in the dataset is the ”Insert” action, which has the highest number of total failures (Figure 4, top right). This is due to the complexity of the insertion motion, which typically involves multiple steps: approaching the target, executing a search pattern to align the object with its socket, and finally applying downward force to insert the object.
Among all insertion tasks, inserting the BNC and bolt 4 are the most difficult. These tasks not only require aligning the object with the socket but also rotating it around the normal axis to the board to secure the object. The higher failure rate for inserting objects like the Ethernet cable or USB cable arises for two reasons: first, both plugs are directional and must have a specific orientation relative to the socket; second, the plugs are located on the edge of Task Board #1, increasing the likelihood of being beyond the robot’s workspace.
The majority of failures in the ”Pick” action (Figure 4, top left) occur because the gripper either misses the object or the object slips out of the gripper while it is closing. Failures in the ”Remove” action (Figure 4, bottom left) often result from improper alignment of the gripper with the object during removal, causing the object to ”jam” in the socket and not be released.
In contrast, we observe the fewest failures for the ”Place” action (Figure 4 bottom right), as it is the simplest to execute successfully. The only requirement is that the object remains on the table after the action. However, failures in this action do occur if the object slips prematurely from the gripper and lands on the task board, which we classify as a failure in this work, as an object on the board might obstruct sockets of other objects.
@@ IV-C Interaction point diversity
Prior works have shown that diversity in the interaction points of actions within a dataset improves generalization and performance on downstream tasks [15]. To ensure high diversity in the collected data, we instructed the operator to randomize the board and object poses for each trial during data collection.
Figure 5 shows the approximate interaction point of all 4,551 demonstrations, obtained by sampling the final end-effector position of each demonstration. The robot is positioned in the origin and faces the positive x-axis direction. In most trials, the objects and the board were placed in front of the robot, within its workspace.
Figure 6: Mean and standard deviation of the force and torque measurements for selected actions. We normalize the action demonstrations by their length to obtain the action’s progress. We can clearly see patterns in the recorded interaction forces and torques. Similar patterns can be found in the remaining actions of the dataset.
@@ IV-D Force and Torque patterns
We analyze whether any patterns occur in the demonstrated force and torque profiles of the actions. Since the demonstrations of each action have different durations, for visualization it is necessary to unify the number of the force and torque measurements to identify patterns across all demonstrations of a specific action. To achieve this, we normalize each demonstration by its duration, converting the demonstrations from the time domain to the ”progress domain.” A progress of 100% indicates the action has finished. Next, we resample each demonstration to a common number of samples (500 in our case) by linearly interpolating the data. Finally, we plot the mean force and torque values along with their standard deviations for each action and visually search for meaningful patterns. The most interesting findings are presented in Figure 6.
Figures 6 A, B, and C show the Z-component of the measured force and the torques around the X and Y axes for the ”Insert waterproof” action. Three distinct phases of the insertion motion are clearly visible: the free-space movement between 0% and 20% of the motion, the first contact and search phase between 20% and 75%, and the final phase starting around 75%. At approximately 75%, the force decreases as the plug aligns with the socket and is released. From this point until the end of the motion, the gripper pushes down on the connector until it snaps into place.
Figures 6 D, E, and F show the Z-component of the measured force and the torques around the Y and Z axes for the ”Insert BNC” action. Here, the different motion phases are also distinguishable. Notably, there is a steady increase in torque around the Z-axis as the plug is tightened until the grooves in the plug reach their limit.
Figures 6 G, H, and I show the Z-component of the measured force and the torques around the Y and Z axes for the ”Place round peg 4” action. The most notable pattern is observed in the Z-force, where the influence of the object’s weight is clearly visible. The force measurement returns to zero after the object is released.
Finally, Figures 6 J, K, and L show the Z-component of the measured force and the torques around the Y and Z axes for the ”Remove small gear” action. As the pulling force along the Z-axis increases, periodic changes in the Z-torque are observed. These changes correspond to the operator pulling on the gear while twisting it back and forth to loosen it.
@@ V-A Temporal Action Segmentation
For the benchmarking purposes, we evaluate the performance of a state-of-the-art visual TAS model, DiffAct [34]. DiffAct leverages diffusion processes [37] to model and learn action boundaries. Diffusion processes work by progressively adding noise to the ground truth information and learning how to iteratively remove this noise. In conditional diffusion processes [38], the denoising process is guided using additional information. For DiffAct, video visual features serve as the conditioning information. We follow the original DiffAct paper and use I3D video features [39].
@@
Figure 7: Temporal action segmentation results. In red, we highlight instances where the ”Pick” action was not predicted by DiffAct. In blue, we mark areas where similar objects were confused. In this case, DiffAct confused ”round peg 1” with ”square peg 1” and ”square peg 2.”
@@
We hypothesize that the lower performance of DiffAct on the REASSEMBLE dataset is due to the increased challenges it presents. Firstly, the REASSEMBLE dataset contains a significantly higher number of unique actions (69 compared to 19 in 50Salads), and many of the objects are less visually distinguishable between each other, such as gears and pegs. Additionally, the REASSEMBLE dataset has almost twice the median number of actions per video (36 compared to 19 in 50Salads). Furthermore, REASSEMBLE often includes sequences where very long actions (e.g., Insert and Remove) are separated by very short actions (e.g., Pick and Place). DiffAct frequently misses these short actions, as illustrated in Figure 7.
We believe the REASSEMBLE dataset presents new opportunities for advancing temporal action segmentation (TAS) in robotics. One underexplored area is the proper fusion of multimodal data in TAS models, which has been limited partly due to a lack of suitable datasets. Moreover, the challenging nature of REASSEMBLE could facilitate research into embedding common-sense reasoning and hierarchical understanding within TAS models. For example, enabling models to learn that before inserting an object, it must first be picked up.
@@ V-B Motion Policy Learning
The primary objective of this study is to introduce a novel robot manipulation dataset specifically designed for contact-rich manipulation tasks, rather than to develop a new policy learning methodology. For the experimental evaluations, we adopt a well-established state-of-the-art approach by integrating Dynamic Movement Primitives (DMPs) into our framework [41]. DMPs have been extensively researched over several decades, providing a robust mathematical foundation for learning and reproducing complex motion trajectories. While contemporary models such as diffusion policies and action chunking transformers have shown promising results across various tasks, they have yet to be optimized for high-precision applications such as assembly tasks involving the NIST board [3, 42]. Consequently, we employ DMPs due to their proven efficacy in precise motion reproduction and adaptability to varying conditions [43]. In our approach, DMPs are utilized to learn the weights corresponding to specific motion conditions, particularly we focus on Cartesian positions. The formulation of the DMPs in our system is governed by the following set of differential equations
@@
Performance demonstrations highlight successful object detection and manipulation tasks, including the challenging assembly and disassembly of gears. As shown in Figure 8, the DMP framework effectively learns the required motion trajectories for these tasks, facilitating precise operations such as picking, inserting, removing, and placing the large gear. For insertion tasks, a spiral motion search is conducted on the $xy$-plane while maintaining a constant downward force along the $z$-axis [44]. This approach utilizes force thresholds learned from the dataset to optimize insertion precision. The spiral search strategy balances the trade-off between compliance and accuracy, ensuring successful gear insertion. By integrating DMPs with our perception pipeline and control strategies, the proposed framework achieves robust, high-precision performance in contact-rich manipulation tasks. We conducted a series of experiments to evaluate the performance of a large gear assembly and disassembly task, encompassing pick, insert, remove, and place motions. Each action was executed 10 times to assess its success rate and failure modes. The pick action was successfully executed in 8 out of 10 trials, with failures occurring due to the gear slipping from the gripper. The insert motion exhibited the lowest success rate, similar to human performance, with 7 out of 10 successful insertions. The primary failure mode for the insertion task occurred when the spiral search trajectory completed, but the gear remained outside the target position. The remove action was successfully performed in 8 out of 10 trials, with failures arising when the gripper mistakenly grasped the plate instead of the gear. This issue was attributed to tracking errors in the impedance controller, which could potentially be mitigated by increasing stiffness. The place motion was successful in all trials. These results highlight the reliability and adaptability of DMPs and force thresholds learned from the dataset in object assembly and disassembly scenarios, demonstrating their effectiveness in executing complex manipulation tasks.
@@ V-C Task Execution and Monitoring using REASSEMBLE
Figure 9 showcases example situations that occur during the execution of the picking task. The leftmost figure represents the initial observation before executing the picking motion. In this case, we observe that ConditionNET predicts the state to be the precondition state of the action, as the gripper is empty. The middle figure shows the situation after the approach motion has been executed and the gear has been grasped. Here, ConditionNET predicts that the effects have been met. The last figure illustrates the case where the gear has fallen out of the gripper. ConditionNET predicts ”Unsatisfied,” as the gripper is closed and no object is being held. For a more detailed overview of the execution monitoring experiments, please refer to the supplementary video.
@@ VI Limitations
Even though we demonstrate that patterns in the torque profiles of each action can be identified, their quality could be improved by using more sophisticated teleoperation devices. The currently employed omega.6 input device provides haptic feedback only in the three translational dimensions, leaving the operator without information about interaction torques. With additional torque feedback, the success rate of insertion and removal tasks could be higher, as the operator would better adapt the orientation of the end-effector to avoid jamming. A viable alternative to the omega.6 is the omega.7 teleoperation device, as it is able to provide torque feedback, though it comes at a substantially higher price. Open-hardware projects such as Gello [46] offer a cheaper alternative; however, they typically rely on less expensive motors, which limit their force and torque reproduction resolution.
Additionally, the currently available action segmentation annotations are provided at only one hierarchical level. These annotations were designed to represent the minimal actions required to describe the assembly and disassembly tasks, thereby minimizing the cognitive load of annotating the data by the teleoperator during data collection. However, each action could be further decomposed into more primitive motions. For example, certain ”Insert” actions could be subdivided into: approach socket $\rightarrow$ search $\rightarrow$ push object into socket. Expanding the dataset with these additional labels would facilitate research into multi-level task hierarchy learning.

### v2

@@ III-D Data Collection
During teleoperation, all sensory information is recorded using the rosbag tool. Subsequently, the raw recordings are post-processed into HDF5 files, whose structure can be found in the Appendix. We do not downsample or synchronize any sensory data during postprocessing; instead, we save the raw data along with timestamps. This approach allows for later adjustments in synchronization or data downsampling for specific downstream tasks. To optimize memory usage, all video and audio recordings are stored as encoded MP4 and MP3 files, respectively, which reduces the dataset size significantly, by a factor of six compared to the unprocessed rosbag recordings.
The accuracy of the predicted text for action narration depends on how clearly the operator spoke during data collection. We observed that some errors occur in the transcriptions; for example, "D-SUB" is often recognized as "the sub." Such errors can be easily corrected automatically by substituting incorrect phrases with the correct ones. Additionally, we manually validate the data stored in the recordings and the action segment and success annotations. For this, we develop and leverage a data visualization tool which we describe in more detail in the Appendix. We verify that all data sources cover the entire recording, are synchronized, and that the action and success labels accurately reflect the events in the videos. If any labelling errors are identified, we correct them. Additionally, if any data is missing, we record the corresponding file name. Finally, based on the known sets of objects and actions, we can compute all valid annotations. We leverage this information to validate whether all action annotations are correct and ensure that no transcription mistakes remain in the dataset.
The skill annotations were manually added after recording all demonstrations. We identified nine commonly shared skills across high-level actions. Grasp refers to free-space movement of the empty gripper until it closes on the object, followed by Lift, which involves moving the object upward. Approach denotes the free-space movement with an object inside of the gripper until the object makes contact with the socket, while Align describes a search motion to correctly position the object with respect to the socket. Release involves letting go of the object, either after alignment or when placing it on the table. Twist is used when rotating a BNC connector or nut, regardless of direction. Push refers to inserting the object into a socket, commonly seen with gears and waterproof connectors, and Pull denotes extracting the object from a socket. Finally, Nudge captures a gentle push to dislodge a peg that may be stuck after release. After completing the annotations, we verified that all skill names were free of typos and manually reviewed infrequent skill-action combinations to ensure they were correctly labeled and not erroneous.
Figure 3: Sankey diagram showing the hierarchical structure and how skills are distributed within actions. The REASSEMBLE dataset contains 121 unique skill-object pairs.
@@ IV-A Action distribution
The number of demonstrations for each action-object pair and their distribution provide insights into the diversity and balance of the REASSEMBLE dataset. To achieve the best performance across all actions, it is crucial that the dataset contains a large and relatively equal number of demonstrations for each action. This ensures that downstream models observe each unique action an equal number of times, reducing performance bias toward more frequently occurring actions. The REASSEMBLE dataset includes four actions and 17 different objects, resulting in a total of 68 unique action-object pairs, or 69 when including the "Idle" action. Figure 4 shows the number of occurrences for each unique action. For clarity, we present one plot per verb in the REASSEMBLE dataset.
The REASSEMBLE dataset contains a minimum of 55 demonstrations for the "Remove square peg 2" action and a maximum of 86 demonstrations for the "Pick USB" and "Insert BNC" actions. This demonstrates that the dataset is balanced.
Figure 3 illustrates the skill composition of each action. The dataset contains a total of 4,551 actions, decomposed into 10,195 skills with an average of 2.2 skills per action. On average, the Pick action consists of three skills, while most other actions typically involve two. The Insert action, which is the most challenging in the REASSEMBLE dataset, shows the greatest variability, with some demonstrations requiring up to ten steps. Insertion sequences with more than three skills often involve objects like the BNC connector or the nut. For instance, screwing the nut requires releasing it, rotating the gripper back, and re-grasping it, due to the limited rotational range of the robot’s final joint. Uncommon action-skill combinations typically occur during failure recovery. For example, a Push during the Place action appears once, when the Ethernet connector gets stuck in the gripper. Similarly, an Align skill appears during Place when a peg needed to be reinserted into its holder.
@@ IV-B Action difficulty and failure modes
The number of failed demonstrations per action can serve as a metric for task difficulty, as operators are more likely to fail when the motion is complex. Figure 5 shows the relationship between successful and unsuccessful demonstrations for each unique action. From the figure, we observe that the most difficult action in the dataset is the "Insert" action, which has the highest number of total failures (Figure 5, top right). This is due to the complexity of the insertion motion, which typically involves multiple steps: approaching the target, executing a search pattern to align the object with its socket, and finally applying downward force to insert the object.
Among all insertion tasks, inserting the BNC and nut 4 are the most difficult. These tasks not only require aligning the object with the socket but also rotating it around the normal axis to the board to secure the object. The higher failure rate for inserting objects like the Ethernet cable or USB cable arises for two reasons: first, both plugs are directional and must have a specific orientation relative to the socket; second, the plugs are located on the edge of Task Board #1, increasing the likelihood of being beyond the robot’s workspace.
The majority of failures in the "Pick" action (Figure 5, top left) occur because the gripper either misses the object or the object slips out of the gripper while it is closing. Failures in the "Remove" action (Figure 5, bottom left) often result from improper alignment of the gripper with the object during removal, causing the object to "jam" in the socket and not be released.
In contrast, we observe the fewest failures for the "Place" action (Figure 5, bottom right). The only requirement is that the object remains on the table after the action. However, failures in this action do occur if the object slips prematurely from the gripper and lands on the task board, which we classify as a failure in this work, as an object on the board might obstruct the sockets of other objects.
@@ IV-C Interaction point diversity
Prior works have shown that diversity in the interaction points of actions within a dataset improves generalization and performance on downstream tasks [17]. To ensure high diversity in the collected data, we instructed the operator to randomize the board and object poses for each trial during data collection. Figure 6 shows the approximate interaction point of all 4,551 demonstrations, obtained by sampling the final end-effector position of each demonstration. The interaction points approximately show the positions of the objects and the board within the robot’s workspace during data collection. The robot is positioned at the origin and faces the positive x-axis direction. In most trials, the objects and the board were placed in front of the robot, within its workspace.
Figure 7: Mean and standard deviation of force and torque measurements for selected actions. To analyze the temporal evolution of each action, we normalize all demonstrations by their duration and compute the average wrench profile over the action progress. Distinct patterns emerge in the force and torque signals across actions, and similar trends are observed in other actions throughout the dataset.
@@ IV-D Force and Torque patterns
We analyze whether any patterns occur in the demonstrated force and torque profiles of the actions. Since the demonstrations of each action have different durations, for visualization it is necessary to unify the number of the force and torque measurements to identify patterns across all demonstrations of a specific action. To achieve this, we normalize each demonstration by its duration, converting the demonstrations from the time domain to the "progress domain." A progress of 100% indicates the action has finished. Next, we resample each demonstration to a common number of samples (500 in our case) by linearly interpolating the data. Finally, we plot the mean force and torque values along with their standard deviations for each action and visually search for meaningful patterns. The most interesting findings are presented in Figure 7.
Figure 7A shows the Z-component of the measured force for the Insert waterproof action. Three distinct phases of the motion are visible: free-space movement from 0% to 20%, initial contact and alignment between 20% and 75%, and the pushing phase starting around 75%. Around this point, the force briefly drops as the plug aligns with the socket and is released. From that moment until the end of the motion, the gripper continues to apply downward force until the connector snaps into place. Figure 7B shows the Z-component of the measured force for the Place round peg 4 action. Here, the effect of the object’s weight is clearly observed. Around the 80% mark, the gripper opens and releases the peg, which results in the force reading returning to nearly zero. Figure 7C shows the torque around the Z-axis for the Remove small gear action. Periodic changes in torque are observed, which correspond to the operator twisting the gear back and forth while pulling it, in order to loosen and remove it.
@@ V-A Temporal Action Segmentation
For benchmarking purposes, we evaluate the performance of a state-of-the-art visual TAS model, DiffAct [37]. DiffAct leverages diffusion processes [40] to model and learn action boundaries. Diffusion processes work by progressively adding noise to the ground truth information and learning how to iteratively remove this noise. In conditional diffusion processes [41], the denoising process is guided using additional information. For DiffAct, video visual features serve as the conditioning information. We follow the original DiffAct paper and use I3D video features [42].
@@
Figure 8: Temporal action segmentation results. In red, we highlight instances where the "Pick" action was not predicted by DiffAct. In blue, we mark areas where similar objects were confused. In this case, DiffAct confused "round peg 1" with "square peg 1" and "square peg 2."
@@
We hypothesize that the lower performance of DiffAct on the REASSEMBLE dataset is due to the increased challenges it presents. Firstly, the REASSEMBLE dataset contains a significantly higher number of unique actions (69 compared to 19 in 50Salads), and many of the objects are less visually distinguishable from each other, such as gears and pegs. Additionally, the REASSEMBLE dataset has almost twice the median number of actions per video (36 compared to 19 in 50Salads). Furthermore, REASSEMBLE often includes sequences where very long actions (e.g., Insert and Remove) are separated by very short actions (e.g., Pick and Place). DiffAct frequently misses these short actions, as illustrated in Figure 8.
One area that remains relatively unexplored is the effective fusion of multimodal data within TAS models, which has been limited partly due to a lack of suitable datasets. In that regard, the REASSEMBLE dataset opens up new possibilities to explore various strategies for integrating multiple data modalities. We further investigated strategies for fusing multimodal data for robotic TAS models [44]. Preliminary results demonstrate improved performance through the integration of visual, auditory, force-torque (wrench), gripper, and pose information. These findings are promising, and we plan to conduct a more comprehensive analysis in future work to further validate and refine our approach.
The REASSEMBLE dataset presents new opportunities for advancing temporal action segmentation (TAS) in robotics. One underexplored area is the proper fusion of multimodal data in TAS models, which has been limited partly due to a lack of suitable datasets. Additionally, the challenging nature of REASSEMBLE could facilitate research into embedding common-sense reasoning and hierarchical understanding within TAS models. For example, enabling models to learn that before inserting an object, it must first be picked up.
@@ V-B Motion Policy Learning
The primary objective of this study is to introduce a novel robot manipulation dataset specifically designed for contact-rich manipulation tasks, rather than to develop a new policy learning methodology. We train and evaluate a language- and goal-conditioned diffusion model for MPL using the REASSEMBLE dataset. The model successfully captures motion profiles for gear assembly and disassembly, achieving a goal-point error of $\approx$1.3 cm, comparable to the SOTA method [45]. While this error is sufficient for tasks like pick and place, it presents challenges for high-precision tasks such as insertion. The contemporary models such as diffusion policies and action chunking transformers have shown promising results across various tasks, they have yet to be optimized for high-precision applications such as assembly tasks involving the NIST board [3, 46]. Therefore, for the complete experimental evaluations, we adopt a well-established state-of-the-art approach by integrating Dynamic Movement Primitives (DMPs) into our framework [47, 48]. DMPs have been extensively researched over several decades, providing a robust mathematical foundation for learning and reproducing complex motion trajectories. In our approach, DMPs are utilized to learn the weights corresponding to specific motion conditions, particularly we focus on Cartesian positions. The formulation of the DMPs in our system is governed by the following set of differential equations
@@
Performance demonstrations highlight successful object detection and manipulation tasks, including the challenging assembly and disassembly of gears. As shown in Figure 9, the DMP framework effectively learns the required motion trajectories for these tasks, facilitating precise operations such as picking, inserting, removing, and placing the large gear. For insertion tasks, a spiral motion search is conducted on the $xy$-plane while maintaining a constant downward force along the $z$-axis [49]. This approach utilizes force thresholds learned from the dataset to optimize insertion precision. The spiral search strategy balances the trade-off between compliance and accuracy, ensuring successful gear insertion. By integrating DMPs with our perception pipeline and control strategies, the proposed framework achieves robust, high-precision performance in contact-rich manipulation tasks. We conducted a series of experiments to evaluate the performance of a large gear assembly and disassembly task, encompassing pick, insert, remove, and place motions. Each action was executed 10 times to assess its success rate and failure modes. The pick action was successfully executed in 8 out of 10 trials, with failures occurring due to the gear slipping from the gripper. The insert motion exhibited the lowest success rate, similar to human performance, with 7 out of 10 successful insertions. The primary failure mode for the insertion task occurred when the spiral search trajectory completed, but the gear remained outside the target position. The remove action was successfully performed in 8 out of 10 trials, with failures arising when the gripper mistakenly grasped the gear holding plate instead of the gear. This issue was attributed to tracking errors in the impedance controller, which could potentially be mitigated by increasing impedance gains. The place motion was successful in all trials. These results highlight the reliability and adaptability of DMPs and force thresholds learned from the dataset in object assembly and disassembly scenarios, demonstrating their effectiveness in executing complex manipulation tasks.
@@ V-C Task Execution and Monitoring using REASSEMBLE
Figure 10 showcases example situations that occur during the execution of the picking task. The leftmost figure represents the initial observation before executing the picking motion. In this case, we observe that ConditionNET predicts the state to be the precondition state of the action, as the gripper is empty. The middle figure shows the situation after the approach motion has been executed and the gear has been grasped. Here, ConditionNET predicts that the effects have been met. The last figure illustrates the case where the gear has fallen out of the gripper. ConditionNET predicts "Unsatisfied," as the gripper is closed and no object is being held. For a more detailed overview of the execution monitoring experiments, please refer to the supplementary video.
@@ VI Limitations
The REASSEMBLE dataset leverages the NIST Task Board#1, which consists of geometrically similar objects with limited visual features. Objects such as round and square pegs and gears pose a particular challenge, as they share similar visual characteristics but differ proprioceptively. This suggests that proprioceptive data may be more effective than visual data for object identification in temporal action segmentation tasks. While this limitation stems from the design of the selected NIST Assembly Task Board, we plan to expand the dataset with additional boards. In particular, Task Board#4 features objects, such as cables, that are more similar in terms of proprioceptive properties but can be distinguished visually. Additionally, the dataset not including depth information constitutes a limitation for robotic perception and manipulation. It was excluded as with the depth data, the size of the dataset would become approximately 12 times larger. Depth data provides three-dimensional spatial context that enables robots to accurately perceive object geometries, estimate distances, and plan precise manipulation trajectories. However, current approaches are developing manipulation methods using only RGB images [3, 29].